OpenStudy (anonymous):

the volume of a sphere is found using the formula V=4/3(3.14)r^3. The area of a circle is found using the formula A=3.14r^2. What is the ratio of the volume to the area?

OpenStudy (anonymous):

dividing the 2, you get (4/3)*r, so that would be the ratio, i believe

OpenStudy (anonymous):

$$\frac{V}{A}=\frac43\frac{\pi r^3}{\pi r^2}=\frac43r$$

OpenStudy (anonymous):

thanks.